The product of the second and third of three consecutive integers is 2 more than 16 times the first integer. Find the integers.
Found 2 solutions by mananth, MathTherapy:
Answer by mananth(16949)

Let the integers be x,x+1,x+2
The product of the second and third of three consecutive integers is 2 more than 16 times the first integer
(x+1)(x+2) = 6x+2
x^2+3x+2=6x+2
x^2-3x=0
x(x-3)=0
x=0 OR x=3
0,1,2
OR
3,4,5
Both satisfy the condition

Answer by MathTherapy(10858)


The product of the second and third of three consecutive integers is 2 more than 16 times the first integer. Find the integers.
Two sets: